how many gallons of a 80% antifreeze solution must be mixed with 100 gallons of 10% antifreeze to get a mixture that is 70% antifreeze
Answer by stanbon(75887)

how many gallons of a 80% antifreeze solution must be mixed with 100 gallons of 10% antifreeze to get a mixture that is 70% antifreeze
-----------------
alcohol + alcohol = alcohol
0.80x + 0.10*100 = 0.70(x+100)
---
80x + 10*100 = 70x + 70*100
10x = 60*100
x = 600 gallons (amt. of 80% solution needed)
----
Cheers,
Stan H.
=================